OpenStudy (anonymous):

The equation of line CD is y = 3x - 3. Write an equation of a line perpendicular to line CD in slope-intercept form that contains point (3, 1). (5 points) y = 3x + 0 y = -3x - 8 y = x + 2 y = - x + 0

OpenStudy (texaschic101):

you need to find the slope of the equation given. We do this because a perpendicular line will have a negative reciprocal slope. Do you know the slope of y = 3x - 3 ??

OpenStudy (anonymous):

no i just think its choice 1

OpenStudy (texaschic101):

no...its not choice 1. I can tell you right now, just by looking, that the answer does not match any of your answer choices. Are you sure they are written correctly ?

OpenStudy (anonymous):

thats what it says on my assignment

OpenStudy (texaschic101):

the slope of the equation given is 3. A perpendicular line will have a negative reciprocal slope. All that means is " flip " the slope and change the sign. So the slope we need is gonna be -1/3....see how I flipped the slope and changed the sign. Now use y = mx + b slope(m) = -1/3 (3,1)...x = 3 and y = 1 now sub y = mx + b 1 = -1/3(3) + b 1 = -1 + b 1 + 1 = b 2 = b your perpendicular line would be : y = -1/3x + 2

OpenStudy (texaschic101):

possibly choice C....if the -1/3 was in there

OpenStudy (anonymous):

thank you.. my school has a lot of errors so i will let them know

OpenStudy (texaschic101):

glad to help :)